A container contains 30 tokens


A container contains 30 green tokens, 15 blue tokens, and 3 red tokens. Two tokens are randomly selected without replacement. Compute P(F|E).

E−you select a blue token first

F−the second token is blue
Answer #1

Total token = 30 + 15 + 3 = 48


15/48 * 15/48 = 0.097656

                        = 9.7656%

answered by: 0bliviant